# A Unified View of Group Fairness Tradeoffs Using Partial Information Decomposition

### v1

@@ IV Tradeoffs Between Unfairness Measures
Conclusion. footnotetext: This work was supported in part by NSF CAREER Award 2340006.By introducing this unifying framework, we provide a tool for gaining a more nuanced understanding of the interplay between different unfairness measures, thereby enhancing the decision-making process in the deployment of fair ML systems. Our work holds broader implications in fields such as algorithmic fairness auditing [18, 6], explainability [55, 56, 57], policy regulation [1], where it can significantly contribute to the evaluation and understanding of unfairness in ML models. This work not only furthers the theoretical discourse but would also have significant societal implications, guiding the trajectory toward more responsible and equitable machine learning in high-stakes settings.
Appendix A Proofs for Theorems in Section III
@@
Appendix B Proofs for Theorems in Section IV

### v2

@@ 6 Conclusion
By introducing this unifying framework, we provide a tool for gaining a more nuanced understanding of the interplay between different unfairness measures, thereby improving the decision-making and deployment of fair machine learning systems. Our work holds broader implications in fields such as fairness auditing [8], explainability [57, 58, 59], policy regulation [1, 14], where it can significantly contribute to the evaluation and understanding of unfairness in machine learning models. This work not only furthers the theoretical discourse but would also have significant societal implications, guiding the trajectory toward more responsible and equitable machine learning in high-stakes settings. Future work could explore efficient methods to estimate PID [60, 61, 49, 62, 50] (also see [45] for more discussion on PID estimation).
Appendix A Proofs for Theorems in Section 3
@@
Appendix B Proofs for Theorems in Section 4